OpenStudy (anonymous):

Find the equation of the line tangent to f(x) = 2/x-1 at the point (2, 2).

OpenStudy (anonymous):

I found the derivative (x-1)^2

OpenStudy (anonymous):

I got the slope 1

OpenStudy (anonymous):

i used y-y1 = m(x-x1)

OpenStudy (anonymous):

I got

hartnn (hartnn):

(2/x) -1 or 2/(x-1) ?

OpenStudy (anonymous):

y=1x

OpenStudy (anonymous):

2/(x-1)

hartnn (hartnn):

then your derivative is incorrect.

OpenStudy (anonymous):

but the options I am given are. a)x+y+3=0

OpenStudy (anonymous):

b)x-y=4 , c) y-2= (x-2)/(x-1)^2, d) y=-2x+6

OpenStudy (anonymous):

or e) none

OpenStudy (anonymous):

why? what did I do wrong?

hartnn (hartnn):

derivative of x^n = nx^(n-1) so, derivative of 1/x is -1/x^2

hartnn (hartnn):

derivative of 2/(x-1) = 2 derivative of (x-1)^(-1)

hartnn (hartnn):

what would be next step ?

OpenStudy (anonymous):

find the derivative? which is: - 1/(x-1)^2

hartnn (hartnn):

yes, so the derivative is -2/(x-1)^2

OpenStudy (anonymous):

so now I insert the 2 in the x?

OpenStudy (anonymous):

hartnn quick question. is +infin. the same as infinity?

hartnn (hartnn):

yes, and yes when we say infinity, its +infinity when we want negative infinity, we must explicitly say -infinity

hartnn (hartnn):

tell me what slope you get when you put x=2

OpenStudy (anonymous):

-2/(2-1)^2 = -2/1 =-2

OpenStudy (anonymous):

thank you

hartnn (hartnn):

yes, you got the equation ?

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

so

OpenStudy (anonymous):

y-2=-2(x-(-)2) y-2 = -2(x+2) y-2 = -2 x -4 y = -2x+6

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you!!!!

hartnn (hartnn):

welcome ^_^